OpenStudy (anonymous):

A ball is dropped from a treetop 256 feet above the ground. How long does it take to hit the ground? (s = 16t^2) 4 seconds 8 seconds 16 seconds

OpenStudy (anonymous):

I think A is the answer...

OpenStudy (cwrw238):

256 = 16t^2 t^2 = 256/16 = 16 t = ?

OpenStudy (cwrw238):

sqrt16

OpenStudy (anonymous):

o_0

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

i think